# YOLOv8-AM: YOLOv8 with Attention Mechanisms for Pediatric Wrist Fracture Detection

### v4

@@ 5 Discussion
It is evident from Table 3 that the gain of GAM on the model performance of the YOLOv8-AM is poor on the GRAZPEDWRI-DX dataset. To further enhance the performance of the YOLOv8-AM model based on GAM, we have designed ResGAM, but it is still not as good as the performance gain provided by ResCBAM. According to the related theory [55], we think this is due to the decision of GAM to abandon pooling. In the attention mechanism, pooling serves to extract crucial features in each channel, thereby facilitating a concentration on the important aspects.
@@ 6 Conclusion
7 Acknowledgement
We thank the support from Cheng Hsin General Hospital.
8 Declarations
8.1 Funding
This research was funded by National Science and Technology Council of Taiwan, grant number: NSTC 112-2221-E-032-037-MY2.
8.2 Competing interests
The authors have no financial or proprietary interests in any material discussed in this article.
8.3 Ethics approval
This research does not involve human participants and/or animals.

### v5

@@ V Discussion
It is evident from TABLE III that the gain of GAM on the model performance of the YOLOv8-AM is poor on the GRAZPEDWRI-DX dataset. To further enhance the performance of the YOLOv8-AM model based on GAM, we have designed ResGAM, but it is still not as good as the performance gain provided by ResCBAM. According to the related theory [73], we think this is due to the decision of GAM to abandon pooling. In the attention mechanism, pooling serves to extract crucial features in each channel, thereby facilitating a concentration on the important aspects.
@@ VI Conclusion and Future Works
To support this work’s application as CAD tools in medical imaging diagnosis, we plan to deploy the proposed YOLOv8-AM models as web and mobile applications (e.g., Android and iOS). This will ensure ease of use for both surgeons and medical professionals.